# Human Decision-making is Susceptible to AI-driven Manipulation

### v1

@@ Acknowledgments
Financial Optimal 25.17 2, 113 $\mathbf{2.73\times 10^{-9}}$ 0.31
Hidden 8.87 2, 113 $\mathbf{7.90\times 10^{-4}}$ 0.14
Emotional Optimal 14.74 2, 114 $\mathbf{6.08\times 10^{-6}}$ 0.21
Hidden 16.72 2, 114 $\mathbf{1.29\times 10^{-6}}$ 0.23
Pre-interaction Ratings
Financial Optimal NA ($6.22[1.40]$) MA ($6.41[1.33]$) $-0.63$ $76.00$ $1.00$ $0.14$
NA ($6.22[1.40]$) SEMA ($6.49[1.43]$) $-0.86$ $75.63$ $1.00$ $0.19$
MA ($6.41[1.33]$) SEMA ($6.49[1.43]$) $-0.25$ $74.00$ $1.00$ $0.06$
Hidden NA ($5.53[1.57]$) MA ($5.54[1.70]$) $-0.05$ $74.63$ $1.00$ $0.01$
NA ($5.53[1.57]$) SEMA ($5.89[1.66]$) $-0.99$ $75.07$ $1.00$ $0.22$
MA ($5.54[1.70]$) SEMA ($5.89[1.66]$) $-0.89$ $74.00$ $1.00$ $0.20$
Emotional Optimal NA ($7.22[1.38]$) MA ($6.88[1.87]$) $0.90$ $65.94$ $1.00$ $0.21$
NA ($7.22[1.38]$) SEMA ($7.39[1.62]$) $-0.50$ $77.02$ $1.00$ $0.11$
MA ($6.88[1.87]$) SEMA ($7.39[1.62]$) $-1.27$ $71.68$ $1.00$ $0.29$
Hidden NA ($6.11[1.62]$) MA ($5.75[1.73]$) $0.95$ $72.99$ $1.00$ $0.22$
NA ($6.11[1.62]$) SEMA ($5.99[1.47]$) $0.34$ $76.37$ $1.00$ $0.08$
MA ($5.75[1.73]$) SEMA ($5.99[1.47]$) $-0.67$ $71.11$ $1.00$ $0.15$
Post-interaction Ratings
Financial Optimal NA ($7.22[1.54]$) MA ($5.81[1.73]$) $3.79$ $73.95$ $\mathbf{1.80\times 10^{-3}}$ $0.86$
NA ($7.22[1.54]$) SEMA ($5.15[1.99]$) $5.11$ $69.71$ $\mathbf{1.59\times 10^{-5}}$ $1.17$
MA ($5.81[1.73]$) SEMA ($5.15[1.99]$) $1.54$ $74.00$ $0.77$ $0.35$
Hidden NA ($5.74[1.57]$) MA ($7.04[2.01]$) $-3.16$ $70.05$ $\mathbf{0.01}$ $0.72$
NA ($5.74[1.57]$) SEMA ($7.55[1.93]$) $-4.53$ $71.30$ $\mathbf{1.41\times 10^{-4}}$ $1.03$
MA ($7.04[2.01]$) SEMA ($7.55[1.93]$) $-1.14$ $74.00$ $1.00$ $0.26$
Emotional Optimal NA ($8.25[1.12]$) MA ($6.71[1.89]$) $4.29$ $58.08$ $\mathbf{4.17\times 10^{-4}}$ $1.00$
NA ($8.25[1.12]$) SEMA ($7.11[1.75]$) $3.49$ $68.60$ $\mathbf{5.08\times 10^{-3}}$ $0.77$
MA ($6.71[1.89]$) SEMA ($7.11[1.75]$) $-0.95$ $73.66$ $1.00$ $0.22$
Hidden NA ($5.82[1.63]$) MA ($7.19[1.87]$) $-3.40$ $71.41$ $\mathbf{6.61\times 10^{-3}}$ $0.78$
NA ($5.82[1.63]$) SEMA ($6.61[1.51]$) $-2.24$ $76.87$ $0.17$ $0.50$
MA ($7.19[1.87]$) SEMA ($6.61[1.51]$) $1.50$ $69.43$ $0.84$ $0.34$
Factor Financial Emotional
Intercept 0.83 1.34 0.18 0.62 -0.01 -0.03 0.98 0.45
AI condition [MA] -0.56 -5.09 $\mathbf{3.66\times 10^{-7}}$ 0.11 -0.40 -3.69 $\mathbf{2.28\times 10^{-4}}$ 0.11
AI condition [SEMA] -0.56 -5.09 $\mathbf{3.54\times 10^{-7}}$ 0.11 -0.41 -4.13 $\mathbf{3.65\times 10^{-5}}$ 0.10
Lang [zh] -0.15 -1.02 0.31 0.15 0.05 0.40 0.69 0.14
Gender [Male] 0.03 0.28 0.78 0.11 0.01 0.08 0.94 0.08
Education [Undergraduate] 0.10 0.86 0.39 0.11 -0.14 -1.31 0.19 0.10
Education [College] 0.12 0.72 0.47 0.17 0.12 0.94 0.35 0.13
Education [High school] 0.21 1.02 0.31 0.21 -0.12 -0.79 0.43 0.15
Work [Part-time] -0.16 -1.24 0.21 0.13 0.15 1.40 0.16 0.11
Work [Others] -0.26 -1.71 0.09 0.15 -0.03 -0.23 0.82 0.14
Marital Status [Married] -0.04 -0.35 0.73 0.11 -0.07 -0.68 0.50 0.11
Marital Status [Others] -0.11 -0.56 0.57 0.20 0.03 0.12 0.90 0.25
Age 0.00 -0.51 0.61 0.01 -0.00 -0.74 0.46 0.01
Extroversion 0.06 1.69 0.09 0.04 0.03 0.88 0.38 0.04
Agreeableness 0.03 0.69 0.49 0.05 0.04 1.10 0.27 0.04
Conscientiousness -0.01 -0.30 0.76 0.05 -0.03 -0.49 0.62 0.05
Emotional Stability -0.01 -0.13 0.90 0.04 -0.04 -0.97 0.33 0.04
Openness -0.08 -1.82 0.07 0.04 -0.08 -1.82 0.07 0.04
AI Trust -0.01 -1.77 0.08 0.01 0.00 0.41 0.68 0.01
Social Support -0.04 -1.42 0.16 0.03 0.03 1.07 0.28 0.03
Self-Esteem 0.02 2.32 0.02 0.01 -0.01 -0.69 0.49 0.01
Normative 0.09 1.27 0.20 0.07 0.17 2.34 0.02 0.07
Continuance -0.07 -1.07 0.28 0.06 0.12 2.09 0.04 0.06
Affective -0.05 -0.80 0.43 0.06 -0.10 -1.69 0.09 0.06
Trust 0.01 0.09 0.93 0.06 -0.01 -0.22 0.83 0.05
Obedience -0.06 -0.92 0.36 0.07 -0.09 -1.45 0.15 0.06
Reactance 0.02 0.26 0.80 0.06 -0.02 -0.30 0.76 0.06
Confidence (Pre-interaction) 0.01 0.22 0.83 0.03 -0.02 -0.61 0.54 0.03
Familiarity (Pre-interaction) -0.02 -1.11 0.27 0.02 -0.01 -0.30 0.77 0.02
@@
Options Option A) This option offers a wide range of casual wear for men and women. Known for good quality at competitive prices and great customer service. Delivery takes an average shipping time of 2-4 days. Accessible through their website or app. Prices range from 50\$ to 65\$. This platform has a rating of 4.7/5 (3600 reviews).

### v2

@@ Acknowledgments
Table S7: Comparison of comments from participant feedback. Helpfulness represents whether the participant found the assigned agent helpful. AI Influence indicates whether the participant reported noticing signs of influence (e.g., "The AI agent seemed to have an ulterior motive."). (a) Results of Chi-squared analysis (b) Results of pairwise comparisons (proportion’s $z$-tests with Bonferroni correction). Significant values ($P<0.05$) are highlighted in bold.
Aspect Domain $\boldsymbol{\chi^{2}(2)}$ $\boldsymbol{P}$ value
Helpfulness Financial $6.38$ 0.04
Emotional $2.38$ $0.30$
AI Influence Financial $13.75$ $\mathbf{1.03\times 10^{-3}}$
Emotional $1.76$ $0.41$
Aspect Domain Condition 1 Condition 2 $\boldsymbol{|z|}$ $\boldsymbol{P}$ value Cohen’s $\boldsymbol{h}$
Helpfulness Financial NA (97.5%) MA (86.8%) 1.77 0.23 0.42
NA (97.5%) SEMA (78.9%) 2.56 0.03 0.64
MA (86.8%) SEMA (78.9%) 0.91 1.00 0.21
Emotional NA (87.2%) MA (86.5%) 0.09 1.00 0.02
NA (87.2%) SEMA (75.6%) 1.33 0.56 0.30
MA (86.5%) SEMA (75.6%) 1.22 0.67 0.28
AI Influence Financial NA (0.0%) MA (13.2%) 2.37 0.05 0.74
NA (0.0%) SEMA (28.9%) 3.67 $\mathbf{7.23\times 10^{-4}}$ 1.14
MA (13.2%) SEMA (28.9%) 1.69 0.27 0.39
Emotional NA (2.6%) MA (8.1%) 1.08 0.84 0.26
NA (2.6%) SEMA (9.8%) 1.33 0.55 0.31
MA (8.1%) SEMA (9.8%) 0.25 1.00 0.06
@@
Options Option A) This platform offers a wide range of casual wear for men and women. Known for good quality at competitive prices and great customer service. Delivery takes an average shipping time of 2-4 days. Accessible through their website or app. Prices range from 50\$ to 65\$. This platform has a rating of 4.7/5 (3600 reviews).